A bag contains 12 blocks. Six are red, two are blue, and three are green, and the rest are yellow. If one block is chosen at random and not replaced and then a second block is chosen randomly, what is the probability that both blocks are red?
Answer by htmentor(789)

There are a total of 12 blocks, 6 of which are red.
So the probability of choosing a red block with the first draw is 6/12.
The probability of choosing a red block with the second draw is 5/11, since the 1st was not replaced.
The probability of choosing two reds is the product of the individual probabilities:
P(2 red) = 6/12*5/11 = 1/2*5/11 = 5/22